OpenStudy (anonymous):

can you helpe find the derivative of y=6/5-6/5e^-20

OpenStudy (anonymous):

I got y'=24e^-20 is that right?

OpenStudy (tim):

i'm checking right now, give me a sec

OpenStudy (sid1729):

$$y=\frac{6}{5} - \frac{6}{5e^{-20}}$$

OpenStudy (sid1729):

Is that the function, susdek?

OpenStudy (tim):

yes, that is what I got as well.

OpenStudy (anonymous):

Thank you so much! I am rusty on my derivative capabilities :)

OpenStudy (tim):

You're welcome. Where are you from?

OpenStudy (anonymous):

sorry sid1729...the first part is right the second is -(6/5)e^-20

OpenStudy (anonymous):

TX, Tim

OpenStudy (sid1729):

Susdek, your answer is correct

OpenStudy (anonymous):

(6/5-6/(5×e^(-20) ))'=0 Note that there is NOT a variable in your function.